# BetterV: Controlled Verilog Generation with Discriminative Guidance

### v1

@@ 4.4.1 Synthesis Nodes Reduction
Besides giving enhancement on functional and syntactic correctness, it is important to consider the hardware attributes. Therefore, we also train the discriminator to improve the performance of Verilog related to later EDA stage, i.e. the synthesis. We make the labels for discriminator depending on whether the logic networks that synthesised from the generated Verilog have less nodes than the reference Verilog. The netlist nodes number are collected by using the “proc; aigmap; stat” commands from $\mathsf{Yosys}$, which transform the Verilog into an And-Inverter-Graph (AIG). The instruction becomes rewriting the reference to have less nodes after synthesis as shown in Figure 4. We select several problems in VerilogEval-human to intuitively demonstrate the performance of BetterV, i.e. the reference nodes number and the generated nodes number are represented. The presented nodes number in the table are the average nodes number of the generated Verilog. As shown in Table 4, with the guidance from discriminator, BetterV can always generate Verilog that has less netlist nodes. The last two columns refer to the proportion of improvements in node reduction compared to BetterV-Base (“Com Base”) and Reference (“Com Ref”) respectively. It can be found that, on average, the Verilog generated by BetterV have $46.52\%$ fewer nodes than the reference model and $31.68\%$ less nodes than BetterV-base.
@@
4.4.2 Verification Runtime Reduction
We make the labels for discriminator depending on whether generated SystemVerilog has less SAT solving time than the reference. The SAT solving time is collected by using the “hierarchy; proc; opt; sat -verify -seq 100 -tempinduct -prove-asserts” commands from $\mathsf{Yosys}$, which solve the SAT problem to prove all the asserts in a circuit with 100 time steps. The instruction becomes rewriting the reference to reduce verification runtime for solving SAT problems, which is the same as the example shown in Figure 4 after replacing part of the description to ”but less Boolean Satisfiability (SAT) solving time”. Since the VerilogEval benchmark doesn’t contain SystemVerilog that includes the assertions inside the design, we choose another benchmark for our experiment in this section. ANSI-C benchmarks give Verilog with safety assertions and can be used for the evaluation of BetterV (Mukherjee et al., 2015). Some designs among them are selected to evaluate the performance of BetterV. As shown in Table 5, after the discrimiative guidance, BetterV is able to further decrease the SAT solving time. We also show the improvement ratio between the verification time of Verilog generated by BetterV and BetterV-base (“Com Base”) and reference Verilog (“Com Ref”) in the last two columns respectively. It shows that the Verilog generated by BetterV can save $22.45\%$ in verification time compared with the reference Verilog, and $13.99\%$ in time compared with BetterV-base. The ability to reduce the verification runtime indicates that LLMs can understand how to rewrite the implementation to accelerate the provement. And it gives the hope that we will be able to solve more complex and complicated problems and hence facilitate the issue of scalability in formal verification.

### v2

@@ 4.3 Customized Generation in BetterV
Besides the measurement on functional correctness, it is important to consider the performance in EDA downstream tasks, which is our customized generation. Note that in the rest of the paper, we conduct the experiments only based on the model CodeLlama-7B-Instruct [22] to demonstrate our results. Therefore, “BetterV” is employed to replace “BetterV-Codellama”, i.e. the model fine-tuned on CodeLlama-7B-Instruct [22]. And “BetterV-base” refers to the base models of “BetterV-CodeLlama”, i.e. the model has undergone instruction tuning but have not yet enhanced by discriminator.
@@ 4.3.1 Synthesis Nodes Reduction
Firstly, we need to consider the hardware attributes. Therefore, we train the discriminator to improve the performance of Verilog related to later EDA stage, i.e. the synthesis. We make the labels for discriminator depending on whether the logic networks that synthesised from the generated Verilog have less nodes than the reference Verilog. The netlist nodes number are collected by using the “proc; aigmap; stat” commands from $\mathsf{Yosys}$, which transform the Verilog into an And-Inverter-Graph (AIG). The instruction becomes rewriting the reference to have less nodes after synthesis as shown in Figure 4. We select several problems in VerilogEval-human to intuitively demonstrate the performance of BetterV, i.e. the reference nodes number and the generated nodes number are represented. The presented nodes number in the table are the average nodes number of the generated Verilog. As shown in Table 2, with the guidance from discriminator, BetterV can always generate Verilog that has less netlist nodes. The last two columns refer to the proportion of improvements in node reduction compared to BetterV-Base (“Com Base”) and Reference (“Com Ref”) respectively. It can be found that, on average, the Verilog generated by BetterV have $46.52\%$ fewer nodes than the reference model and $31.68\%$ less nodes than BetterV-base.
@@
4.3.2 Verification Runtime Reduction
We make the labels for discriminator depending on whether generated SystemVerilog has less SAT solving time than the reference. The SAT solving time is collected by using the “hierarchy; proc; opt; sat -verify -seq 100 -tempinduct -prove-asserts” commands from $\mathsf{Yosys}$, which solve the SAT problem to prove all the asserts in a circuit with 100 time steps. The instruction becomes rewriting the reference to reduce verification runtime for solving SAT problems, which is the same as the example shown in Figure 4 after replacing part of the description to ”but less Boolean Satisfiability (SAT) solving time”. Since the VerilogEval benchmark doesn’t contain SystemVerilog that includes the assertions inside the design, we choose another benchmark for our experiment in this section. ANSI-C benchmarks give Verilog with safety assertions and can be used for the evaluation of BetterV [18]. Some designs among them are selected to evaluate the performance of BetterV. Note that for the “Rotate” case, we uncomment all the assertions to enhance the difficulty. As shown in Table 3, after the discrimiative guidance, BetterV is able to further decrease the SAT solving time. We also show the improvement ratio between the verification time of Verilog generated by BetterV and BetterV-base (“Com Base”) and reference Verilog (“Com Ref”) in the last two columns respectively. It shows that the Verilog generated by BetterV can save $22.45\%$ in verification time compared with the reference Verilog, and $13.99\%$ in time compared with BetterV-base. The ability to reduce the verification runtime indicates that LLMs can understand how to rewrite the implementation to accelerate the provement. And it gives the hope that we will be able to solve more complex and complicated problems and hence facilitate the issue of scalability in formal verification.
4.4 Abltion Study
4.4.1 Impact of Discriminator
We also show that the generative discriminator has the ability to further enhance the performance after the guidance. In the task of functional correctness, the labels of discriminator are made by checking whether the generated modules are functional equivalence with the reference module. The equivalence checking is done by the $\mathsf{eqy}$ tool in $\mathsf{Yosys}$. With the problems in VerilogEval-human, the results in Table 4 illustrate that the guidance from the discriminator can not only enhance the capability on our fine-tuned LLMs, i.e. these base models, but also on the original pre-trained LLMs, i.e. CodeLlama. This observation indicates that our trained discriminator can be employed to any LLM, only if they have the same vocabulary size.
Table 4: Impact of discriminator on functional correctness.
Model pass@1 pass@5 pass@10
CodeLlama 18.2 22.7 24.3
CodeLlama + Dis 20.3 24.1 24.7
BetterV-CodeLlama-base 40.0 49.5 53.0
BetterV-CodeLlama 40.9 50.0 53.3
Table 5: Impact of discriminator on syntactic correctness.
Model pass@1 pass@5 pass@10
CodeLlama 41.5 50.8 53.8
CodeLlama + Dis 49.1 58.4 61.1
BetterV-CodeLlama-base 82.6 97.6 99.2
BetterV-CodeLlama 87.1 98.2 99.3
5 Discussion
5.1 Current limitations
We have shown that BetterV can achieve remarkable performance on Verilog generation and the generative discriminator enables improvements on Verilog-related EDA downstream tasks. While BetterV contributes valuable insights and a practical method for enhancing the performance of open-source LLMs, it still has limitations on closed-source models due to the need for access to token-level probabilities. The research about guidance generation on closed-source models remains an unsolved issue.
In addition, while the generative discriminator uses only a small model, the additional computational overhead it brings cannot be ignored. Research to reduce this overhead is needed to accelerate the guidance in an accuracy-optimal manner. For example, it is possible to use model inference acceleration methods such as quantization and pruning on the discriminator.
5.2 Applicability to other domains
The potential applicability of our methods to other domains with similar challenges of data scarcity and precision requirements is promising. For example, many domains utilize structured languages (e.g., SQL for databases, LaTeX for typesetting, and XML for web services). The principles behind BetterV’s instruct-tuning and generative discriminators can be adapted to improve the generation of such languages, particularly where precision and correctness are non-negotiable. Fields such as genomics, where data is highly structured and errors can have significant consequences, could benefit from our approach. The instruct-tuning methodology could be adapted to understand the syntax and semantics of genetic sequences, while generative discriminators could be used to guide the generation of sequences to complex biological constraints.